What is the solution of 2(x+3)^1/3-5=1?: What is the solution of 2(x+3)^1/3-5=1?
Answer by jim_thompson5910(9897)

2(x+3)^(1/3)-5=1 Start with the given equation


2(x+3)^(1/3)=6 Add 5 to both sides.


(x+3)^(1/3)=3 Divide both sides by 2


x+3=27 Cube both sides


x=24 Subtract 3 from both sides.